# FineCLIPER: Multi-modal Fine-grained CLIP for Dynamic Facial Expression Recognition with AdaptERs

### v1

@@ Abstract.
Dynamic Facial Expression Recognition (DFER) is crucial for understanding human behavior. However, current methods exhibit limited performance mainly due to the scarcity of high-quality data, the insufficient utilization of facial dynamics, and the ambiguity of expression semantics, etc. To this end, we propose a novel framework, named Multi-modal Fine-grained CLIP for Dynamic Facial Expression Recognition with AdaptERs (FineCLIPER), incorporating the following novel designs: 1) To better distinguish between similar facial expressions, we extend the class labels to textual descriptions from both positive and negative aspects, and obtain supervision by calculating the cross-modal similarity based on the CLIP model; 2) Our FineCLIPER adopts a hierarchical manner to effectively mine useful cues from DFE videos. Specifically, besides directly embedding video frames as input (low semantic level), we propose to extract the face segmentation masks and landmarks based on each frame (middle semantic level) and utilize the Multi-modal Large Language Model (MLLM) to further generate detailed descriptions of facial changes across frames with designed prompts (high semantic level). Additionally, we also adopt Parameter-Efficient Fine-Tuning (PEFT) to enable efficient adaptation of large pre-trained models (i.e., CLIP) for this task. Our FineCLIPER achieves SOTA performance on the DFEW, FERV39k, and MAFW datasets in both supervised and zero-shot settings with few tunable parameters. Analysis and ablation studies further validate its effectiveness. Project page: https://haroldchen19.github.io/FineCLIPER-Page/
@@
copyright: noneconference: ; 2024journalyear: ;booktitle: ACM, Conferenceccs: Computing methodologies Computer visionccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
Facial expressions are important signals to convey human emotions, thus accurately recognizing them has significant meaning for various tasks, including interpersonal communication, human-computer interaction (HCI) (Liu et al., 2017; Ladak et al., 2024; Hundt et al., 2024), mental health diagnosing (Kumar et al., 2024; Sarkar et al., 2023; Bisogni et al., 2022), driving safety monitoring (Wilhelm, 2019; Wang et al., 2023b; Yan et al., 2024a), etc. Traditional Facial Expression Recognition (FER) resorts to static images. However, since dynamic emotional changes could not be well-represented within a single image, research attention has been shifted to Dynamic Facial Expression Recognition (DFER), which distinguishes the temporally displayed facial expressions in videos.
The study of DFER algorithms starts from highly-controlled environments (Cao et al., 2014; Livingstone and Russo, 2018a; Martin et al., 2006), where faces are frontal and non-blurry (Li and Deng, 2022; Wang et al., 2022a) as shown in Fig. 1 (a). However, such an ideal assumption makes the obtained models vulnerable to real-world situations. Therefore, researchers have turned to more open scenes and constructed several in-the-wild DFER datasets, e.g., DFEW (Jiang et al., 2020), FERV39k (Wang et al., 2022b), and MAFW (Liu et al., 2022a), to facilitate the development of corresponding methods (Zhao and Liu, 2021a; Wang et al., 2023a; Li et al., 2023c; Ma et al., 2022; Li et al., 2022; Liu et al., 2023) as demonstrated in Fig. 1 (b). While category labels are treated without semantic meanings (e.g., Happiness may only be represented by a class id ”0”), recent research on DFER (Zhao and Patras, 2023; Foteinopoulou and Patras, 2024; Tao et al., 2024; Li et al., 2023b) has further delved into the exploration of vision-language multi-modal learning beyond the traditional classification paradigm based on CNNs (Kossaifi et al., 2020; Sun et al., 2020; Fan et al., 2016; Lee et al., 2019), RNNs (Baddar and Ro, 2019; Wang et al., 2022c), and transformers (Li et al., 2023c; Zhao and Liu, 2021b).
Figure 1. Frameworks for Dynamic Facial Expression Recognition.
@@
Furthermore, we adopt a semantically hierarchical strategy to comprehensively mine useful information from the input video data. Specifically, features from directly embedding video frames stand at a relatively low semantic level. For middle semantic level, we utilize a well-trained face analysis model (i.e. FaceXFormer (Narayan et al., 2024)) to extract the face segmentation masks and landmarks from each frame. Intuitively, the former offers prior about face structures while the latter provides specific pivots for model attention. Additionally, we try to obtain descriptions at a high semantic level for describing dynamic facial changes across frames. This is realized by leveraging a well-trained MLLM, Video-LLaVA (Lin et al., 2023) to act as a facial expression analyst following given template-based prompts, and the generated descriptions will be carefully refined. All the above features at various semantic levels will be integrated to obtain the final representation of a given video.
To summarize, our contributions are as follows:
  • We introduce FineCLIPER, a novel multi-modal framework that enhances Dynamic Facial Expression Recognition (DFER) through extensively mining useful information at different semantic levels from the video data, and all the obtained features (i.e., features embedded from visual frames, face segmentation, face landmarks, and the extra fine-grained descriptions obtained via MLLM) are integrated finally to serve as a more comprehensive overall representation;
@@ 2. Related Work
Dynamic Facial Expression Recognition. In early DFER research, the focus was on developing diverse local descriptors on lab-controlled datasets (Cao et al., 2014; Livingstone and Russo, 2018b, a). Then the rise of deep learning and accessible in-the-wild DFER datasets (Jiang et al., 2020; Wang et al., 2022b; Liu et al., 2022a) leads to new trends towards DFER research. The first trend (Fan et al., 2016; Kossaifi et al., 2020; Lee et al., 2019) involves the direct use of 3D CNNs (Tran et al., 2015; Hara et al., 2018; Tran et al., 2018) to extract joint spatio-temporal features from raw videos. The second trend (Ebrahimi Kahou et al., 2015; Kollias and Zafeiriou, 2020; Sun et al., 2020; Wang et al., 2022c) combines 2D CNNs (Simonyan and Zisserman, 2014; Chung et al., 2014) with RNNs (Chung et al., 2014; Hochreiter and Schmidhuber, 1997) for feature extraction and sequence modeling. The third emerging trend integrates transformer (Dosovitskiy et al., 2020), as demonstrated in works like Former-DFER (Zhao and Liu, 2021b), STT (Ma et al., 2022), and IAL (Li et al., 2023c). These methods combine convolutional and attention-based approaches to enhance the understanding of visual data, especially in distinguishing samples based on varying visual dynamics. However, in prior efforts, the semantic meaning of class labels is neglected, and insufficient attention has been paid to the subtle and nuanced movements of the human face. Therefore, based on the well-trained large cross-modal models (i.e. CLIP), we propose to extend the class label to textual supervision both positively and negatively. Moreover, to fully exploit the visual information within videos, we also design a hierarchical information mining strategy to generate representative video features, which is a weighted fusion of various features involving different semantic levels, including video frame feature, the middle-level facial semantics from segmentation maps and detected landmarks, we well as the high-level semantics encoded from fine-grained descriptions provided by MLLM.
@@
CLIP in Classification. With the development of computer vision (He et al., 2016; Chen et al., 2024; Huang et al., 2024), Vision-Language Models (VLMs), e.g., CLIP (Radford et al., 2021), have recently demonstrated superior performance across various tasks (Wang et al., 2024; Ni et al., 2022; Yu et al., 2024; Deng et al., 2023; Hyung et al., 2023; Huang et al., 2023; Liu et al., 2024; Yan et al., 2024b; Teng et al., 2021). CLIP leverages a vast corpus of image-text pairs to ground its framework in contrastive learning, resulting in robust pre-trained image and text encoders that demonstrate remarkable feature extraction capabilities. Recent studies (Zhao and Patras, 2023; Tao et al., 2024; Li et al., 2023b) have also applied CLIP to the DFER task. Among them, A3lign-DFER (Tao et al., 2024) introduces a comprehensive alignment paradigm for DFER through a complicated design. CLIPER (Li et al., 2023b) adopts a two-stage training paradigm instead of end-to-end training; however, it is limited in capturing temporal information. Furthermore, DFER-CLIP (Zhao and Patras, 2023) incorporates a transformer-based module to better capture temporal information in videos, but it requires fully fine-tune the image encoder and the proposed temporal module during training, leading to inefficiency.
@@ 3. Methodology
In this section, we first briefly go through the overall pipeline and basic notations of the framework in Sec. 3.1. Then, we elaborate on how to augment the original class labels to obtain positive-negative textual supervision in Sec. 3.2, followed by details about our hierarchical information mining strategy to obtain multi-modal features in Sec. 3.3. The integration of diverse features is introduced in Sec. 3.4. The overall pipeline is illustrated in Fig. 2.
@@ 3.3. Hierarchical Information Mining
The average number of tokens in our refined descriptions is approximately 35 tokens. However, research (Zhang et al., 2024) demonstrates the actual effective length of CLIP’s text encoder is even less than 20 tokens. Hence, to better explore the fine-grained description of facial changes, we adopt the text encoder of Long-CLIP (Zhang et al., 2024) as our fine-grained text encoder $\mathbf{E}_{F}$, which can support text inputs of up to 248 tokens. The refined fine-grained description, denoted as $F$, is further tokenized and projected into embeddings $\mathbf{X}_{F}$. Following a procedure similar to the text encoder described in Sec. 3.2, the input is further constructed as $\mathbf{z}_{F}^{(0)}=\mathbf{X}_{F}+\mathbf{E}_{F}$, where $\mathbf{E}_{F}$ is the positional encoding of $F$. Subsequently, by feeding it into the projector $\mathbf{h}_{F}$, we could contain the final feature vector of $F$ as: $\mathbf{f}=\mathbf{h}_{F}(\mathbf{z}_{F,l}^{(L_{F})})$.
@@
and the final similarity is obtained by: $sim_{i,\mathbf{m}}=sim_{i,\mathbf{m}}^{pos}-sim_{i,\mathbf{m}}^{neg}$, which further distinguishes similarity among similar categories
@@ 4.2. Main Results
Figure 6. Attention visualizations for DFEW w.r.t. two ground-truth expression labels—’Happiness’ (Top) and ’Surprise’ (Bottom).
4.3. Ablation Studies

### v2

@@ Abstract.
Dynamic Facial Expression Recognition (DFER) is crucial for understanding human behavior. However, current methods exhibit limited performance mainly due to the insufficient utilization of facial dynamics, and the ambiguity of expression semantics, etc. To this end, we propose a novel framework, named Multi-modal Fine-grained CLIP for DFER with AdaptERs (FineCLIPER), incorporating the following novel designs: 1) To better distinguish between similar facial expressions, we extend the class labels to textual descriptions from both positive and negative aspects, and obtain supervision by calculating the cross-modal similarity based on the CLIP model; 2) Our FineCLIPER adopts a hierarchical manner to effectively mine useful cues from DFE videos. Specifically, besides directly embedding video frames as input (low semantic level), we propose to extract the face segmentation masks and landmarks based on each frame (middle semantic level) and utilize the Multi-modal Large Language Model (MLLM) to further generate detailed descriptions of facial changes across frames with designed prompts (high semantic level). Additionally, we also adopt Parameter-Efficient Fine-Tuning (PEFT) to enable efficient adaptation of large pre-trained models (i.e., CLIP) for this task. Our FineCLIPER achieves SOTA performance on the DFEW, FERV39k, and MAFW datasets in both supervised and zero-shot settings with few tunable parameters. Project page: https://haroldchen19.github.io/FineCLIPER-Page/
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28–November 1, 2024; Melbourne, VIC, Australia.booktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28–November 1, 2024, Melbourne, VIC, Australiaisbn: 979-8-4007-0686-8/24/10doi: 10.1145/3664647.3680827ccs: Computing methodologies Computer visionccs: Human-centered computing Human computer interaction (HCI)
Figure 1. Frameworks for DFER.
@@ 1. Introduction
Facial expressions are important signals to convey human emotions, thus accurately recognizing them has significant meaning for various tasks, including interpersonal communication, human-computer interaction (HCI) (Liu et al., 2017; Ladak et al., 2024; Hundt et al., 2024), mental health diagnosing (Kumar et al., 2024; Sarkar et al., 2023; Bisogni et al., 2022), driving safety monitoring (Wang et al., 2023b; Yan et al., 2024a; Zhang et al., 2024b), etc. Traditional Facial Expression Recognition (FER) resorts to static images. However, since dynamic emotional changes could not be well-represented within a single image, research attention has been shifted to Dynamic Facial Expression Recognition (DFER), which distinguishes the temporally displayed facial expressions in videos.
The study of DFER algorithms starts from highly-controlled environments (Cao et al., 2014; Livingstone and Russo, 2018a; Martin et al., 2006), where faces are frontal and non-blurry (Li and Deng, 2022; Wang et al., 2022a) as shown in Fig. 1 (a). However, such an ideal assumption makes the obtained models vulnerable to real-world situations. Therefore, researchers have turned to more open scenes and constructed several in-the-wild DFER datasets, e.g., DFEW (Jiang et al., 2020), FERV39k (Wang et al., 2022b), and MAFW (Liu et al., 2022a), to facilitate the development of corresponding methods (Zhao and Liu, 2021a; Wang et al., 2023a; Li et al., 2023c; Ma et al., 2022; Li et al., 2022; Liu et al., 2023) as demonstrated in Fig. 1 (b). While category labels are treated without semantic meanings (e.g., Happiness may only be represented by a class id ”0”), recent research on DFER (Zhao and Patras, 2023; Foteinopoulou and Patras, 2024; Tao et al., 2024; Li et al., 2023b) has further delved into the exploration of vision-language multi-modal learning beyond the traditional classification paradigms (Kossaifi et al., 2020; Sun et al., 2020; Baddar and Ro, 2019; Wang et al., 2022c; Li et al., 2023c; Zhao and Liu, 2021b).
@@
Furthermore, we adopt a semantically hierarchical strategy to comprehensively mine useful information from the input video data. Specifically, features from directly embedding video frames stand at a relatively low semantic level. For middle semantic level, we utilize a well-trained face analysis model (i.e. FaceXFormer (Narayan et al., 2024)) to extract the face segmentation masks and landmarks from each frame. Intuitively, the former offers prior about face structures while the latter provides specific pivots for model attention. Additionally, we try to obtain descriptions at a high semantic level for describing dynamic facial changes across frames. This is realized by leveraging a well-trained MLLM, Video-LLaVA (Lin et al., 2023) to act as a facial expression analyst following given template-based prompts, and the generated descriptions will be carefully refined. All the above features at various semantic levels will be integrated to obtain the final representation of a given video. To summarize, our contributions are as follows:
  • [leftmargin=*]
  • We introduce FineCLIPER, a novel multi-modal framework that enhances Dynamic Facial Expression Recognition (DFER) through extensively mining useful information at different semantic levels from the video data, and all the obtained features (i.e., features embedded from visual frames, face segmentation, face landmarks, and the extra fine-grained descriptions obtained via MLLM are integrated to serve as a more comprehensive representation;
@@ 2. Related Work
Dynamic Facial Expression Recognition. In early DFER research, the focus was on developing diverse local descriptors on lab-controlled datasets (Cao et al., 2014; Livingstone and Russo, 2018b, a). Then the rise of deep learning and accessible in-the-wild DFER datasets (Jiang et al., 2020; Wang et al., 2022b; Liu et al., 2022a) leads to new trends towards DFER research. The first trend (Fan et al., 2016; Kossaifi et al., 2020; Lee et al., 2019) involves the direct use of 3D CNNs (Tran et al., 2015; Hara et al., 2018; Tran et al., 2018) to extract joint spatio-temporal features from raw videos. The second trend (Ebrahimi Kahou et al., 2015; Kollias and Zafeiriou, 2020; Sun et al., 2020; Wang et al., 2022c) combines 2D CNNs (Simonyan and Zisserman, 2014; Chung et al., 2014) with RNNs (Chung et al., 2014; Hochreiter and Schmidhuber, 1997) for feature extraction and sequence modeling. The third emerging trend (Zhao and Liu, 2021b; Ma et al., 2022; Li et al., 2023c) integrates transformer (Dosovitskiy et al., 2020). These methods combine convolutional and attention-based approaches to enhance the understanding of visual data, especially in distinguishing samples based on varying visual dynamics. However, in prior efforts, the semantic meaning of class labels is neglected, and insufficient attention has been paid to the subtle and nuanced movements of the human face. Therefore, based on the well-trained large cross-modal models (i.e., CLIP), we propose to extend the class label to textual supervision both positively and negatively. Moreover, to fully exploit the visual information within videos, we also design a hierarchical information mining strategy to generate representative video features, which is a weighted fusion of various features involving different semantic levels, including video frame feature, the middle-level facial semantics from segmentation maps and detected landmarks, as well as the high-level semantics encoded from fine-grained descriptions.
@@
CLIP in Classification. With the development of computer vision (Chen et al., 2024; Huang et al., 2024; Shao et al., 2020a, 2018, b; Wang et al., 2023c), Vision-Language Models (Jiang et al., 2024), e.g., CLIP (Radford et al., 2021), have recently demonstrated superior performance across various tasks (Wang et al., 2024; Yu et al., 2024; Yan et al., 2024b). Recent studies (Zhao and Patras, 2023; Tao et al., 2024; Li et al., 2023b) have also applied CLIP to the DFER task. Among them, A3lign-DFER (Tao et al., 2024) introduces a comprehensive alignment paradigm for DFER through a complicated design. CLIPER (Li et al., 2023b) adopts a two-stage training paradigm instead of end-to-end training; however, it is limited in capturing temporal information. Furthermore, DFER-CLIP (Zhao and Patras, 2023) incorporates a transformer-based module to better capture temporal information in videos, but it requires fully fine-tune the image encoder and the proposed temporal module during training, leading to inefficiency.
@@ 3. Methodology
This section outlines the framework’s overall pipeline and basic notations (Sec. 3.1), the augmentation of class labels for positive-negative textual supervision (Sec. 3.2), the hierarchical information mining strategy (Sec. 3.3), and the integration of these features (Sec. 3.4). The complete pipeline is shown in Fig. 2.
@@ 3.3. Hierarchical Information Mining
The average number of tokens in our refined descriptions is approximately 35 tokens. However, research (Zhang et al., 2024a) demonstrates the actual effective length of CLIP’s text encoder is even less than 20 tokens. Hence, to better explore the fine-grained description of facial changes, we adopt the text encoder of Long-CLIP (Zhang et al., 2024a) as our fine-grained text encoder $\mathbf{E}_{F}$, which can support text inputs of up to 248 tokens. The refined fine-grained description, denoted as $F$, is further tokenized and projected into embeddings $\mathbf{X}_{F}$. Following a procedure similar to the text encoder described in Sec. 3.2, the input is further constructed as $\mathbf{z}_{F}^{(0)}=\mathbf{X}_{F}+\mathbf{E}_{F}$, where $\mathbf{E}_{F}$ is the positional encoding of $F$. Subsequently, by feeding it into the projector $\mathbf{h}_{F}$, we could contain the final feature vector of $F$ as: $\mathbf{f}=\mathbf{h}_{F}(\mathbf{z}_{F,l}^{(L_{F})})$.
@@
is obtained by: $sim_{i,\mathbf{m}}=sim_{i,\mathbf{m}}^{pos}-sim_{i,\mathbf{m}}^{neg}$, which further distinguishes similarity among similar categories
@@ 4.3. Ablation Studies
Figure 6. Attention visualizations for DFEW w.r.t. two ground-truth expression labels—‘Happiness’ (Top) and ‘Surprise’ (Bottom).
@@ 5. Conclusion
Acknowledgements.
This work is mainly supported by the National Natural Science Foundation of China (NSFC) under Grant 62306239. This work is also supported by the Young Talent Fund of Association for Science and Technology in Shaanxi, China.